the length of a rectangle is 4cm more than its width. if the area of the rectangle is 60cm^2 find the dimensions.
Answer by mananth(16946)

let the width be x
length = x+4
Area = L*W
x(x+4)=60
x^2+4x=60
x^2+4x-60=0
x^2+10x-6x-60=0
x(x+10)-6(x+10)=0
(x+10)(x-6)=0
choose the positive value since length cannot be negative
x=6 the width
Length =10cm